OpenStudy (hockeychick23):

A scientist is working on a new test for cotton allergies. She knows that 2.5% of all the students at Annie's school are allergic to cotton. For those who are allergic, 80% test positive and 20% test negative. For those who are not allergic, the test shows 10% positive and 90% negative.Use Bayes's theorem to find the probability that a student at Annie's school is allergic to cotton, given that the test results are positive. Bayes Theorem: P(A l B)= [P(B I A) x P(A)]/ P(B)

OpenStudy (ganpat):

I am not very sure about this answer.. but worth a try.. lets someone else confirm it.. Bayes Theorem: P(A l B)= [P(B I A) x P(A)]/ P(B) 1. Test on the students who she knew are allergic.. p(A) = 0.025 p(B) = 0.2 p(B/A) = 0.8 so, p(A/B) = [0.8 *0.025]/0.2 = 0.1 2. similarly find for students who she thought are not allergic and test was run on them.. 3. Add the two figures..

OpenStudy (mathmate):

Yes, you could use Bayes theorem: \(P(A|B)=\large \frac{P(B | A) \times P(A)}{P(B)}\) Let events A=allergic, A'=not alergic B=tests positive, B'=tests negative and use Bayes Theorem to calculate P(A|B), or Probability that a student is allergic to cotton, given the test results are positive. From the given data, 2.5% are allergic, so \(P(A)=0.025\) Out of the allergic, 80% test positive, i.e. 80% test positive GIVEN that they are allergic. so \(P(B|A)=0.80\). The last quantity we need is P(B), which is not directly given. We will need the law of total probability to find P(B). \(P(B)=P(B\cap \Omega)=P(B\cap (A+A'))=P(B\cap A)+P(B\cap A')\) where \(P(B\cap A)=P(B|A)*P(A)=0.8*0.025\) test positive and alergic \(P(B\cap A')=P(B|A')*P(A')=0.1*0.975\) test positive and not alergic using \(P(X|Y)=P(X\cap Y)/P(Y)\) Putting it altogether you'd get the required probability P(B).

OpenStudy (mathmate):

An easier way is to draw a contingency diagram (at least to check your results) from which we deduce right away that P(A|B)=0.02/0.11175=0.1702. However, the question requires you to use Baye's theorem, so you still have to go through the motions.